# Torsion of $\mathbb{Q}$-curves over number fields of small odd prime degree

### v1

@@ 3.5. Eliminating $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$
Proof.   Suppose that this group does occur for some $\mathbb{Q}$-curve $E$ over a number field $K$. We first use Lemma 2.7 with $M=2$, $N=6$ to conclude that there is a $2$-isogenous curve $E_{1}$ such that $\rho_{E_{1},24}(G_{K})$ consists (in a suitable basis) of matrices of the form $\begin{bmatrix}h&*\\ 0&*\end{bmatrix}$, where $h\in\{1,13\}$. The curve $E_{1}$ is also $K$-isogenous to an elliptic curve $E^{\prime}/K$ with a rational $j$-invariant. As before, we may assume that this isogeny is cyclic.
Now, using Proposition 2.5 with $H=\{1,13\}$, we obtain that there is an $\alpha$ dividing $24$ such that $\rho_{E^{\prime},24}(G_{K})$ is conjugate to a subgroup of $\textnormal{GL}_{2}(\mathbb{Z}/24\mathbb{Z})$ of the form
$\begin{bmatrix}h&\frac{24}{\alpha}u\\ \alpha v&z\end{bmatrix},$
with $h\in\{1,13\}$.
Twisting to an elliptic curve $E^{\prime\prime}$ defined over $\mathbb{Q}$, we see that $\rho_{E^{\prime\prime},24}(G_{K})$ is conjugate to a subgroup of $\textnormal{GL}_{2}(\mathbb{Z}/24\mathbb{Z})$ of the form
$\begin{bmatrix}h&\frac{24}{\alpha}u\\ \alpha v&z\end{bmatrix},$
with $h\in\{\pm 1,\pm 13\}\leq\mathbb{Z}/24\mathbb{Z}$.
If $\alpha\in\{1,3,8,24\}$, then $E^{\prime\prime}$ has a $24$-isogeny over $K$. However, $24$ is not a possible cyclic isogeny degree of an elliptic curve with a rational $j$-invariant over a cubic field by [17].
In the remaining cases when $\alpha\in\{2,4,6,12\}$, $E^{\prime\prime}$ certainly has a $3$-isogeny over $K$, and it also has a $2$-isogeny and a $4$-isogeny which are independent. This $3$-isogeny is rational by Proposition 2.8. Suppose that the $4$-isogeny is not defined over $\mathbb{Q}$. Then by [17, Proposition 3.2], the image $\rho_{E^{\prime\prime},4}(G_{\mathbb{Q}})$ is a specific group $H_{20}$ from the list in [18]. However, from the proof of [17, Lemma 3.11], it follows that there is no $3$-isogeny over $\mathbb{Q}$ in this case, a contradiction.
Thus, $E^{\prime\prime}$ has a $4$-isogeny over $\mathbb{Q}$. Note that the $2$-isogeny which is independent from it also has to be defined over $\mathbb{Q}$, as the full $2$-torsion of an elliptic curve with a rational point of order $2$ is defined either over $\mathbb{Q}$ or over a quadratic extension. This means that $E^{\prime\prime}$ has a rational $12$-isogeny and a $2$-isogeny, and they are independent.
We can now use Lemma 2.7 to find a rational elliptic curve with a $24$-isogeny, which is a contradiction, since such elliptic curves don’t exist. $\Box$
4. Torsion over number fields of degree $5$ or $7$
@@ 4.1. Possible prime power orders of points
Let $E$ be a $\mathbb{Q}$-curve over a number field $K$ of degree $p\geq 5$. Suppose that $E$ has a point of order $\ell\neq p$, where $\ell$ is an odd prime.
Then there is an elliptic curve $E_{1}$ which is $K$-isogenous to $E$ such that $E_{1}$ has rational $j$-invariant and a point of order $\ell$.
Proof.   Let $E$ be a $\mathbb{Q}$-curve over $K$. Let $E^{\prime}$ be the $K$-isogenous elliptic curve with rational $j$-invariant. We can factor the isogeny $E\to E^{\prime}$ by writing it down as $E\to E_{1}\to E^{\prime}$, where the isogeny $E\to E_{1}$ has order coprime to $\ell$ and $E^{\prime}$ has order $\ell^{k}$ for some $k\geq 0$. Then $E_{1}$ is a $\mathbb{Q}$-curve isogenous to $E$ and it also has a point of order $\ell$ over $K$.
If $\ell^{k}=1$, then $E_{1}$ is isomorphic to $E$ and hence has rational $j$-invariant.
Otherwise, let $E^{\prime\prime}$ be the quadratic twist of $E$ which is defined over $\mathbb{Q}$. Furthermore, let $E^{\prime\prime\prime}$ be the corresponding quadratic twist of $E_{1}$, such that we have the following diagram (the vertical lines are quadratic twists, and the horizontal lines are $\ell^{k}$-isogenies).
[…]
By [17, Lemma 4.4. and Lemma 4.7.], it follows that the isogeny $E^{\prime\prime}\to E^{\prime\prime\prime}$, from the diagram, which is a priori defined over $K$, is in fact already defined over $\mathbb{Q}$. In particular, $E^{\prime\prime\prime}$ is also defined over $\mathbb{Q}$, so $E_{1}$ has rational $j$-invariant.
$\Box$
@@
In particular, by Theorem 1.5 there are no points of prime order greater than $11$ over degree $5$ extensions, and by [16, Proposition 7.7], there are no points of prime order greater than $7$ over degree $7$ extensions.
In fact, any $\mathbb{Q}$-curve with a point of order $11$ is isogenous to an elliptic curve with $j$-invariant in $\{-11^{2},-11\cdot 131^{3}\}$. Since those curves have no isogenies of odd prime order and no $2$-isogenies over quintic extensions, their isogeny class over quintic extensions does not increase in size, so they are the only $\mathbb{Q}$-curves over quintic extensions with a point of order $11$.
Thus, we may now only focus on powers of $2,3,5,7$.
@@
Proof.
Note that $5$ and $7$ do not divide the orders of $\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})$ and $\textnormal{GL}_{2}(\mathbb{Z}/3\mathbb{Z})$.
Now suppose that there is a $\mathbb{Q}$-curve $E$ over a degree $5$ or $7$ extension with a $16$-isogeny or a $27$-isogeny and let $E^{\prime}$ be an isogenous curve with rational $j$-invariant, $E^{\prime\prime}$ the quadratic twist of $E^{\prime}$ which is rational and $E^{\prime\prime\prime}$ the corresponding twist of $E$.
Then, we may without loss of generality assume that the isogeny $E\to E^{\prime}$ is a $2^{a}3^{b}$-isogeny as other isogeny degrees preserve the $2$-torsion and the $3$-torsion.
However, then the corresponding $2^{a}3^{b}$-isogeny $E^{\prime\prime}\to E^{\prime\prime\prime}$ is defined over $\mathbb{Q}$, so $E^{\prime\prime\prime}$ is a rational elliptic curve and $E$ has rational $j$-invariant. We are now done by Theorem 1.5 and Theorem 1.6.
$\Box$
@@
Proof.   This follows immediately from [8, Theorem 2.], since a Sylow pro-$5$ subgroup’s index is not divisible by $25$Box
@@
Proof.   Suppose that $E$ is a $\mathbb{Q}$-curve over a degree $5$ extension with a point of order $125$, and let $E^{\prime}$ be the $K$-isogenous elliptic curve with rational $j$-invariant, and $E^{\prime\prime}/\mathbb{Q}$ the twist of $E^{\prime}$ which is rational. Then, by Corollary 2.6, $\rho_{E^{\prime\prime},125}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form
$\begin{bmatrix}*&5^{a}\cdot x\\ 5^{3-a}\cdot y&\pm 1\end{bmatrix}$
@@
This means that $\rho_{E^{\prime\prime},125}(G_{K})$ has at most $2\cdot 25\cdot 4\cdot 125=2^{3}\cdot 5^{5}$ elements, so the index is divisible $3\cdot 2^{2}\cdot 5^{4}$, and the index of $\rho_{E^{\prime\prime},125}(G_{\mathbb{Q}})$ is divisible by $3\cdot 2^{2}\cdot 5^{3}$.
By Proposition 2.8, $E^{\prime\prime}$ has a $5$-isogeny over $\mathbb{Q}$. By Lemma 4.3, the index of $\rho_{E^{\prime\prime},5^{\infty}}(G_{\mathbb{Q}})$ is not divisible by $25$, so we reach a contradiction.
Similarly, if $E$ is defined over a degree $7$ extension $K$ and has a point of order $25$, $\rho_{E^{\prime\prime},25}(G_{K})$ is contained in the subgroup of the group of all matrices of the form
$\begin{bmatrix}*&5^{a}\cdot x\\ 5^{2-a}\cdot y&\pm 1\end{bmatrix}$
for some $a\in\{0,1,2\}$. Since $7\nmid|\textnormal{GL}_{2}(\mathbb{Z}/5\mathbb{Z})|$, it follows that $\rho_{E^{\prime\prime},25}(G_{K})=\rho_{E^{\prime\prime},25}(G_{\mathbb{Q}})$. We can now compute that the index of $\rho_{E^{\prime\prime},25}(G_{\mathbb{Q}})$ is at least $3\cdot 2^{2}\cdot 5^{2}$, which is again divisible by $25$, a contradiction with Lemma 4.3. $\Box$
@@
Proof.   Let $E$ be a $\mathbb{Q}$-curve over a degree $5$ extension $K$ with a point of order $49$, $E^{\prime}$ be the $K$-isogenous curve with rational $j$-invariant and $E^{\prime\prime}$ the twist of $E^{\prime}$ which is rational. Then by Corollary 2.6, $E^{\prime\prime}$ has two independent $7$-isogenies over $K$, and since $5\nmid|\textnormal{GL}_{2}(\mathbb{Z}/7\mathbb{Z})|$, they are also defined over $\mathbb{Q}$. However, there then exists a $\mathbb{Q}$-isogenous elliptic curve with a $49$-isogeny over $\mathbb{Q}$, which is a contradiction.
@@
Now let $E$ be a $\mathbb{Q}$-curve over a degree $7$ extension $K$ with a point of order $49$, $E^{\prime}$ be the $K$-isogenous curve with rational $j$-invariant and $E^{\prime\prime}$ the twist of $E^{\prime}$ which is rational. Again, by Corollary 2.6, it follows that $\rho_{E^{\prime\prime},49}(G_{K})$ is a subgroup of the group of all matrices of the form
$\begin{bmatrix}*&7^{a}\cdot x\\ 7^{2-a}\cdot y&\pm 1\end{bmatrix}$
@@
for some $a\in\{0,1,2\}$. In particular, it has a $7$-isogeny over $K$. By Proposition 2.8, it also has a $7$-isogeny over $\mathbb{Q}$. By [13, Theorem 3.9], it follows that the $7$-adic representation over $\mathbb{Q}$ is defined modulo $7$, i.e. it contains all matrices congruent to $I$ modulo $7$.
@@
To summarize, $\rho_{E^{\prime\prime},49}(G_{\mathbb{Q}})$, once a suitable basis has been chosen, has a subgroup of index $7$ contained in the group of matrices of the form $\begin{bmatrix}*&7^{a}\cdot x\\ 7^{2-a}\cdot y&\pm 1\end{bmatrix}$, which has order $7^{3}\cdot 2\cdot 6$, and contains the group of all matrices congruent to $I$ modulo $7$, which has order $7^{4}$. This means that the order of $\rho_{E^{\prime\prime},49}(G_{K})$ must be divisible by $7^{3}$.
@@
Without loss of generality, we may assume that $\rho_{E^{\prime\prime},49}(G_{K})$ also contains $-I$. Furthermore, $\rho_{E^{\prime\prime}.49}(G_{K})$ must have surjective determinant in $(\mathbb{Z}/7\mathbb{Z})^{\times}$. Since fibers of all elements of $(\mathbb{Z}/7\mathbb{Z})\times$ have the same number of elements, and this number is divisible by $2$ (since $\det(-I)=1$), it follows that $\rho_{E^{\prime\prime},49}(G_{K})$ has order divisible by $12$. But then the order of $\rho_{E^{\prime\prime},49}(G_{K})$ is $7^{3}\cdot 12$, so it equals the group of all matrices of the form $\begin{bmatrix}*&7^{a}\cdot x\\ 7^{2-a}\cdot y&\pm 1\end{bmatrix}$.
Then $\rho_{E^{\prime\prime},49}(G_{\mathbb{Q}})$ contains all matrices of the form $\begin{bmatrix}*&7^{a}\cdot x\\ 7^{2-a}\cdot y&\pm 1+7z\end{bmatrix}$ as it contains all matrices congruent to the identity mod $7$, and this is the entire image. We now see that $E^{\prime\prime}$ either has a $49$-isogeny or two independent $7$-isogenies over $\mathbb{Q}$, which is impossible. $\Box$

### v2

@@ 3.5. Eliminating $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$
Proof.   Suppose that this group does occur for some $\mathbb{Q}$-curve $E$ over a number field $K$. We first use Lemma˜2.7 with $M=2$, $N=6$ to conclude that there is a $2$-isogenous curve $E_{1}$ such that $\rho_{E_{1},24}(G_{K})$ consists (in a suitable basis) of matrices of the form $\begin{bmatrix}h&*\\ 0&*\end{bmatrix}$, where $h\in\{1,13\}$. In particular, $E_{1}$ has a $K$-rational $24$-isogeny. The curve $E_{1}$ is also $K$-isogenous to an elliptic curve $E^{\prime}$ with a rational $j$-invariant. As before, we may assume that this isogeny is cyclic of degree $2^{a}\cdot 3^{b}$ for $a,b\geq 0$. Let us factor this isogeny as $E_{1}\to E_{2}$ of degree $3^{b}$ and $E_{2}\to E^{\prime}$ of degree $2^{a}$. If $E_{2}$ has rational $j$-invariant, then $E_{2}$ has a $3$-isogeny and an $8$-isogeny over $K$, so it has a $24$-isogeny as well. However, this is not a valid isogeny degree over cubic fields by [16, Theorem 1.2]. Thus, $E_{2}$ does not have rational $j$-invariant and hence $a>0$. If $a>2$, then $E^{\prime}$ has an $8$-isogeny and a $3$-isogeny, which is again impossible for the same reason. We know that $E_{2}$ has an $8$-isogeny coming from the $8$-isogeny of $E_{1}$. Consider the path in the $2$-isogeny graph formed by these isogenies. Since the distance between $E^{\prime}$ and $E_{2}$ in the graph is $1$ or $2$, the position of $E_{2}$ with respect to this path is among those marked by $\star$ in the following diagram (note that the diagram does not show the entire $2$-isogeny graph).
$E_{2}$$\star_{3}$$\star_{5}$$\bullet$$\star_{1}$$\star_{4}$$\star_{2}$}}
If $E^{\prime}$ is in the position of $\star_{1}$, $\star_{2}$ or $\star_{4}$, it has an $8$-isogeny, a contradiction. In other cases, $E^{\prime}$ has a $4$-isogeny to an elliptic curve with non-rational $j$-invariant and an independent $2$-isogeny, so $\rho_{E^{\prime},2}(G_{K})$ is trivial. Since $E^{\prime}$ has a $4$-isogeny to an elliptic curve with non-rational $j$-invariant, it follows from [16, Proposition 3.2] that some quadratic twist $E^{\prime\prime}$ of $E^{\prime}$ which is defined over $\mathbb{Q}$ has $2$-adic Galois image $\rho_{E^{\prime\prime},2^{\infty}}(G_{\mathbb{Q}})$ equal to the group $H_{20}$ from the database of Rouse and Zureick–Brown in [17]. Reducing modulo $4$, this is a group of order $12$ generated by $\begin{bmatrix}1&1\\ 3&0\end{bmatrix}$ and $\begin{bmatrix}3&3\\ 0&1\end{bmatrix}$. The group $\rho_{E^{\prime\prime},4}(G_{K})$ is a subgroup of order $4$ which is contained inside the group of uppertriangular matrices. There is only one such subgroup. However, if this subgroup is equal to $\rho_{E^{\prime\prime},4}(G_{K})$, then $\rho_{E^{\prime\prime},2}(G_{K})$ is not trivial, because $\rho_{E^{\prime\prime},4}(G_{K})$ contains $\begin{bmatrix}3&3\\ 0&1\end{bmatrix}$. This is a contradiction since then $E^{\prime\prime}$ (and hence also $E^{\prime}$) does not have all $2$-isogenies defined over $K$. $\Box$
4. Torsion over number fields of degree $5$ or 77
@@ 4.1. Possible prime power orders of points
Let $E$ be a $\mathbb{Q}$-curve over a number field $K$ of prime degree $p\geq 5$. Let $\ell\neq p$ be an odd prime. Suppose that $E$ has a point of order $\ell\neq p$ defined over $K$.
Then there exists an elliptic curve $E_{1}$ which is $K$-isogenous to $E$ such that $E_{1}$ has rational $j$-invariant and a point of order $\ell$.
Proof.   Let $E$ be a $\mathbb{Q}$-curve over $K$. Let $E^{\prime}$ be the $K$-isogenous elliptic curve with rational $j$-invariant. We can factor the isogeny $E\to E^{\prime}$ by writing it down as $E\to E_{1}\to E^{\prime}$, where the isogeny $E\to E_{1}$ has degree coprime to $\ell$ and the isogeny $E_{1}\to E^{\prime}$ has order $\ell^{k}$ for some $k\geq 0$. Then $E_{1}$ is a $\mathbb{Q}$-curve isogenous to $E$ which also has a point of order $\ell$ over $K$.
If $\ell^{k}=1$, then $E_{1}$ is isomorphic to $E$ and hence has rational $j$-invariant. Otherwise, let $E^{\prime\prime}$ be the quadratic twist of $E^{\prime}$ which is defined over $\mathbb{Q}$. Furthermore, let $E^{\prime\prime\prime}$ be the corresponding quadratic twist of $E_{1}$, such that we have the following diagram (the vertical lines are quadratic twists, and the horizontal lines are $\ell^{k}$-isogenies).
$E_{1}$$E^{\prime}$$E^{\prime\prime\prime}$$E^{\prime\prime}$$\ell^{k}$$\ell^{k}$twisttwist
Note that $p$ does not divide $\ell\cdot(\ell-1)$ unless $p=5$ and $\ell=11$, so in those cases $\rho_{E^{\prime\prime},\ell^{k}}(G_{\mathbb{Q}})=\rho_{E^{\prime\prime},\ell^{k}}(G_{K})$ and it follows that any $\ell^{k}$-isogeny of $E^{\prime\prime}$ defined over $K$ is in fact already defined over $\mathbb{Q}$.
Now suppose that $\ell=11$ and $p=5$. Then $E^{\prime\prime}$ is a rational elliptic curve with an $11$-isogeny over a quintic extension, so by Proposition˜2.8 it must have a rational $11$-isogeny. However, then $j(E^{\prime\prime})=j(E^{\prime})$ belongs to the set $\{-11^{2},-11\cdot 131\}$. The elliptic curves with these $j$-invariants have only one $11$-isogeny (the rational one) over quintic extensions of $\mathbb{Q}$. They have no isogenies of odd prime order over $\mathbb{Q}$, so their isogeny class does not increase when taking quintic extensions. It follows that they are the only $\mathbb{Q}$-curves over quintic extensions with a point of order $11$. $\Box$
@@
In particular, by Theorem˜1.6 there are no points of prime order greater than $11$ over degree $5$ extensions, and by [15, Proposition 7.7], there are no points of prime order greater than $7$ over degree $7$ extensions. Since the points of order $11$ have just been discussed, we may now only focus on powers of $2,3,5,7$.
@@
Proof.  Note that $5$ and $7$ do not divide the orders of $\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})$ and $\textnormal{GL}_{2}(\mathbb{Z}/3\mathbb{Z})$.
Suppose that there is a $\mathbb{Q}$-curve $E$ over a degree $5$ or $7$ extension with a $16$-isogeny or a $27$-isogeny and let $E^{\prime}$ be an isogenous curve with rational $j$-invariant, $E^{\prime\prime}$ the quadratic twist of $E^{\prime}$ which is rational and $E^{\prime\prime\prime}$ the corresponding twist of $E$, as in the diagram in the proof of Lemma˜4.1. Without loss of generality we may assume that the isogeny $E\to E^{\prime}$ is a $2^{a}3^{b}$-isogeny for some $a,b\geq 0$.
However, the corresponding $2^{a}3^{b}$-isogeny $E^{\prime\prime}\to E^{\prime\prime\prime}$ is defined over $\mathbb{Q}$ since $5$ and $7$ do not divide $\textnormal{GL}_{2}(\mathbb{Z}/2\mathbb{Z})$ or $\textnormal{GL}_{2}(Z/3\mathbb{Z})$, so $E^{\prime\prime\prime}$ is a rational elliptic curve and its twist $E$ has rational $j$-invariant. The claim now follows by Theorem˜1.6 and Theorem˜1.7. $\Box$
@@
Proof.   This follows immediately from [8, Theorem 2.], since the index of a Sylow pro-$5$ subgroup of $\textnormal{GL}_{2}(\mathbb{Z}_{5})$ is not divisible by $25$. $\Box$
@@
Proof.   Suppose that $E$ is a $\mathbb{Q}$-curve defined over a quintic extension of $\mathbb{Q}$ with a point of order $125$. We know that $E$ is $K$-isogenous to an elliptic curve with rational $j$-invariant $E^{\prime}$. Without loss of generality, we may assume that the isogeny degree is $5^{a}$ for some $a\geq 1$. If $a>3$, then $E^{\prime}$ has a $625$-isogeny over a quintic extension, which is impossible by [16, Theorem 1.3]. Thus $a\leq 3$. The elliptic curve $E$ also has a $125$-isogeny defined over $K$ whose kernel is the point of order $125$ on $E$. Since the $5$-isogeny subgraph is a line by Corollary˜2.11, it follows that the isogeny $E\to E^{\prime}$ either factors through the $125$-isogeny whose kernel is the torsion point or the isogenies are independent. If they are independent, then $E^{\prime}$ has a $5^{3+a}$-isogeny, which is again a contradiction. Thus, $E\to E^{\prime}$ factors through the $125$-isogeny. Then, by Corollary˜2.6, $\rho_{E^{\prime},125}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form
$\begin{bmatrix}1&5^{a}\cdot x\\ 5^{3-a}\cdot y&*\end{bmatrix}$
@@
If $E^{\prime\prime}$ is a quadratic twist of $E^{\prime}$ which is a rational elliptic curve, then $\rho_{E^{\prime\prime},125}(G_{K})$ is conjugate to a subgroup of the group of all matrices of the form
$\begin{bmatrix}\pm 1&5^{a}\cdot x\\ 5^{3-a}\cdot y&*\end{bmatrix}$
for some $a\in\{0,1,2,3\}$.
This means that $\rho_{E^{\prime\prime},125}(G_{K})$ has at most $2\cdot 25\cdot 4\cdot 125=2^{3}\cdot 5^{5}$ elements. It follows that the index of $\rho_{E^{\prime\prime},125}(G_{K})$ is divisible by $\textnormal{GL}_{2}(\mathbb{Z}/125\mathbb{Z})/(2^{3}\cdot 5^{5})=2^{2}\cdot 3\cdot 5^{4}$, so it is divisible by $5^{4}$. It follows that the index of $\rho_{E^{\prime\prime},125}(G_{\mathbb{Q}})$ is divisible by $5^{3}$. Since $E^{\prime\prime}$ has a $5$-isogeny over $\mathbb{Q}$, Lemma˜4.3 implies that $\rho_{E^{\prime\prime},5^{\infty}}(G_{\mathbb{Q}})$ is not divisible by $25$, so we reach a contradiction.
Now suppose that $E$ is defined over a degree $7$ extension $K$ and has a point of order $25$. Let $E^{\prime}$ be the isogenous curve which has rational $j$-invariant and let $5^{a}$ be the isogeny degree. We have $a\leq 2$ since $125$ is not a possible isogeny degree of an elliptic curve with rational $j$-invariant.
The $5$-isogeny graph is a line so we can again conclude that the isogeny $E\to E^{\prime}$ either factors through the isogeny whose kernel is a point of order $25$ or they are independent. If they are independent, then $E^{\prime}$ has a $125$-isogeny over an extension of degree $7$, a contradiction, so $E\to E^{\prime}$ factors through the isogeny whose kernel is the torsion point. If we let $E^{\prime\prime}$ be the quadratic twist of $E^{\prime}$ which is defined over $\mathbb{Q}$, we know by Corollary˜2.6 that $\rho_{E^{\prime\prime},25}(G_{K})$ is contained in the subgroup of the group of all matrices of the form
$\begin{bmatrix}\pm 1&5^{a}\cdot x\\ 5^{2-a}\cdot y&*\end{bmatrix}$
for some $a\in\{0,1,2\}$. Since $7\nmid|\textnormal{GL}_{2}(\mathbb{Z}/5\mathbb{Z})|$, it follows that $\rho_{E^{\prime\prime},25}(G_{K})=\rho_{E^{\prime\prime},25}(G_{\mathbb{Q}})$. We can now compute that the index of $\rho_{E^{\prime\prime},25}(G_{\mathbb{Q}})$ is at least $3\cdot 2^{2}\cdot 5^{2}$, which is again divisible by $25$, a contradiction with Lemma˜4.3. $\Box$
@@
Proof.   Let $E$ be a $\mathbb{Q}$-curve over a degree $5$ extension $K$ with a point of order $49$, $E^{\prime}$ be the $K$-isogenous curve with rational $j$-invariant and $E^{\prime\prime}$ the twist of $E^{\prime}$ which is rational. We may assume that the isogeny $E\to E^{\prime}$ has degree $7^{a}$ for $a\geq 1$. Since $7^{2}$ is not a possible isogeny degree of elliptic curves with rational $j$-invariant over quintic fields, it follows that $a=1$ and that the isogeny $E\to E^{\prime}$ factors through the isogeny whose kernel is the point of order $49$. By Proposition˜2.5, $E^{\prime}$ has two independent $7$-isogenies over $K$. Then $E^{\prime\prime}$ also has two independent $7$-isogenies. Since $5\nmid|\textnormal{GL}_{2}(\mathbb{Z}/7\mathbb{Z})|$, they are also defined over $\mathbb{Q}$. However, there then exists a $\mathbb{Q}$-isogenous elliptic curve with a $49$-isogeny over $\mathbb{Q}$, which is a contradiction.
@@
Now suppose that $E$ is a $\mathbb{Q}$-curve over a degree $7$ extension $K$ with a point of order $49$, $E^{\prime}$ the $K$-isogenous curve with rational $j$-invariant and $E^{\prime\prime}$ the twist of $E^{\prime}$ which is rational. We may assume that the isogeny $E\to E^{\prime}$ has degree $7^{a}$ for $a\geq 1$. Since $7^{3}$ is not a possible isogeny degree of an elliptic curve with rational $j$-invariant over a septic field, it follows that the isogeny $E\to E^{\prime}$ factors through the isogeny whose kernel is the point of order $49$. Furthermore, since a twist $E^{\prime\prime}$ of $E^{\prime}$ has a rational $7$-isogeny, we may assume that we have the following diagram.
$E{{E^{\prime}}}$$E_{2}$}}
Here $E^{\prime}$ and $E_{2}$ both have rational $j$-invariant.
by Proposition˜2.5, it follows that for the quadratic twist $E^{\prime\prime}$ of $E^{\prime}$ the group $\rho_{E^{\prime\prime},49}(G_{K})$ is a subgroup of the group of all matrices of the form
$\begin{bmatrix}\pm 1&7\cdot x\\ 7\cdot y&*\end{bmatrix}$
@@
Since $E^{\prime\prime}$ has a $7$-isogeny over $\mathbb{Q}$, [13, Theorem 3.9] implies that that the $7$-adic representation of $E^{\prime\prime}$ over $\mathbb{Q}$ is defined modulo $7$, i.e. it contains all matrices congruent to $I$ modulo $7$.
@@
To summarize, $\rho_{E^{\prime\prime},49}(G_{\mathbb{Q}})$, once a suitable basis has been chosen, has a subgroup of index $7$ contained in the group of all matrices of the form $\begin{bmatrix}\pm 1&7\cdot x\\ 7\cdot y&*\end{bmatrix}$, which has order $7^{3}\cdot 2\cdot 6$ and $\rho_{E^{\prime\prime},49}(G_{\mathbb{Q}})$ contains the group of all matrices congruent to $I$ modulo $7$, which has order $7^{4}$. This means that the order of $\rho_{E^{\prime\prime},49}(G_{K})$ must be divisible by $7^{3}$.
@@
Without loss of generality, taking another quadratic twist if needed, we may assume that $\rho_{E^{\prime\prime},49}(G_{K})$ also contains I-I. We know that $\rho_{E^{\prime\prime}.49}(G_{K})$ has surjective determinant in $(\mathbb{Z}/7\mathbb{Z})^{\times}$. The preimages under $\det:\rho_{E^{\prime\prime},7}(G_{K})\to(\mathbb{Z}/7\mathbb{Z})^{\times}$ of all elements of $(\mathbb{Z}/7\mathbb{Z})\times$ have the same number of elements, and this number is divisible by $2$ since $\det(-I)=1$. It follows that $\rho_{E^{\prime\prime},7}(G_{K})$ and consequently $\rho_{E^{\prime\prime},49}(G_{K})$, has order divisible by $12$. But then $\rho_{E^{\prime\prime},49}(G_{K})$ is divisible by $7^{3}\cdot 12$, and $7\cdot\#\rho_{E^{\prime\prime},49}(G_{K})=\#\rho_{E^{\prime\prime},49}(G_{\mathbb{Q}})$ divides $7^{4}\cdot 12$. It follows that we must have equality everywhere, so $\rho_{E^{\prime\prime},49}(G_{K})$ is equal to the group of all matrices of the form $\begin{bmatrix}\pm 1&7\cdot x\\ 7\cdot y&*\end{bmatrix}$. Since $\rho_{E^{\prime\prime},49}(G_{\mathbb{Q}})$ contains all matrices of this form and all matrices congruent to $I\pmod{7}$, it follows that $\rho_{E^{\prime\prime},49}(G_{\mathbb{Q}})$ is the group of all matrices of the form $\begin{bmatrix}\pm 1+7\cdot z&7\cdot x\\ 7\cdot y&*\end{bmatrix}$. But then $E^{\prime\prime}$ has two independent rational $7$-isogenies, a contradiction. $\Box$